OpenStudy (lefinnois):

for a right circular cone, if the radius is reduced to1/5 its original size and the slant height is reduced to 1/6 its original size, find the surface area if the original radius is 8 centimeters and the original slant height is 13 centimeters.

OpenStudy (lefinnois):

Is it 18.93?

OpenStudy (mrnood):

No slant area = pi *r *l original r = 8 but this is reduces to 1/5 ot the original original l = 13 but this is reduced to 1/6 of teh original so work out the new r and l then calculate pi*r*l

OpenStudy (michele_laino):

hint: using data from the text of the problem, we can write that the new area \(S'\), is a function of the original area \(S\), namely the area computed from the non-reduced, or original dimensions, so we have: \[S' = \pi r'h' = \pi \cdot \frac{r}{5} \cdot \frac{h}{6} = \frac{{\pi rh}}{{30}} = \frac{S}{{30}}\]

OpenStudy (lefinnois):

Working on it! :)

OpenStudy (mrnood):

Both methods are identical One converts the dimensions before calculating the area The other converts the area after calcularting original pi * r * l is all you need for this

OpenStudy (lefinnois):

But that would make it 326.56 and that wouldn't be the answer..

OpenStudy (lefinnois):

Oh, sorry, forgot the reduction part. :P Never mind.

OpenStudy (michele_laino):

hint: please compute this: \(S=\pi \times 8 \times 13=...?\)

OpenStudy (lefinnois):

Yes, that's326.56. Just trying to convert the 8 and 13.

OpenStudy (lefinnois):

Would it be 18.84?

OpenStudy (mrnood):

No

OpenStudy (mrnood):

As has been shown above: take the ORIGINAL area pi * r * l and divide by 30

OpenStudy (lefinnois):

Ah, 18.63

OpenStudy (mrnood):

No - I don't think that is correct either show me pi *r * l with original values

OpenStudy (lefinnois):

Ah, so it's 18.74.

OpenStudy (lefinnois):

You agree @MrNood ?

OpenStudy (mrnood):

no please show how you worked this out step by step

OpenStudy (mrnood):

OK wait - I may be mistaken I read it as asking for the slant area - but it asks for the 'surface area' My mistake - wait a moment the formula is A = pi*r *l + p*r^2 = pi * r (r+l)

OpenStudy (mrnood):

18.93 is correct I apologise for your wasted time....